Calculate the energy of the (n=1) to (n=2) transition of the H atom in electron volts, wave numbers, wavelength and frequency. (this is college upper level Physical Chemistry)
Answer #1

DE = 13.6(1/n1^2 - 1/n2^2)

   = 13.6((1/1^2)-(1/2^2))

   = 10.2 ev

(1/L) = 109677(1/n1^2 - 1/n2^2)

1/L = 109677((1/1^2) - (1/2^2))

L = wavelength = 1.21*10^-7 m

wave number = 1/L = 8.26*10^6 m-1

frequency = 3*10^8/(1.21*10^-7) = 2.5*10^15 s-1